OpenStudy (anonymous):

solve for x sqrt (x) +3 = 1/2 (x+3)

OpenStudy (anonymous):

\[\sqrt{x} + 3 = 1/2 (x+3)\]

OpenStudy (anonymous):

dont see how that works.

OpenStudy (anonymous):

where did the sqrt of x go from the initial problem?

OpenStudy (anonymous):

i can multiply by 2 leaving 2sqrt (x) +6 = x +3 then subtract 3 leaving 2sqrt (x) +3 = x

OpenStudy (anonymous):

do i then square both sides to get 4x+9=x^2

OpenStudy (anonymous):

crap... sorry i coppied the problem wrong... the answer is x=9

OpenStudy (anonymous):

okay going back to 2sqrt(x)+3=x solve for x

OpenStudy (anonymous):

how does x =9?

OpenStudy (anonymous):